# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v5

@@ 4.2 Numerical Results
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks. Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
@@
Figure 4: Comparison between standard Kaiming initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
@@
4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 5: Approximation produced with standard methodologies (shown before ReLU is applied). The neuron colors here are arbitrary; they do not correspond to Figure 2.
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Occasionally, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Effect of Reparameterized Initialization vs Pretraining
Figure 6: Pretraining (right) results in better structural retention in deeper layers.
Figure 6 compares the top performing reparameterized models for $x^{3}$ with and without pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure around layer 4 or 5, devolving into noisy patterns and resulting in higher errors. Pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure during its optimization in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training. This phenomenon highlights the need for theoretical work to derive more expressive training controls, which can give better guidance for the optimization of functions of greater dimension or complexity.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
Figure 7: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
@@
The results presented so far are limited to one-dimensional convex functions, but these are not limitations of our method. Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 9 in the Appendix for a schematic). For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs. Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
@@
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising. In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world datasets and problems with up to eight input dimensions.
@@
We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1. This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957; 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions. But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees. Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
@@
Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
@@ Appendix A Appendix
A.1 Initialization and Training Algorithm
@@
     A-Gradient $\leftarrow\text{Derivative}(\text{Network},A)$triangleright Backpropagate through weight setting
@@
     $A\leftarrow A-\epsilon\times\text{Gradient}$triangleright Update A, Not the network
@@ A.3 Mathematical Results
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. This odd split around layer $i$ makes the proofs cleaner.
@@
Figure 10 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next. Valleys in one layer remain valleys in all future layers since 0 is a fixed point of each $T_{i}$. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition. On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected. Due to the reflection, the slopes of $W_{i}$ on each side of a peak or valley are proportional. Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$. Either way, it’s important to note that the slopes on each side of a bend scale identically during each composition.
@@ A.10 Real-World Data and Classification Problems
Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 15 and 16 already show impressive results using our current approach. In the case of the concrete problem (Figure 15), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced. In the case of the two-spiral classification task (Figure 16), our network is able to learn an accurate decision boundary (left subfigure), whereas a standard network constrained to be block diagonal fails to learn (right subfigure). When dense weight matrices are trained, our networks are almost able to tie with the standard initializations, although convergence can take longer. These results suggest that our method might be quite powerful when its parameter budget is spent on greater width and depth of convex function blocks (as in Figure 9) instead of filling in dense matrices.
@@
A.11 Large-Scale Image Classification Examples
@@
We evaluated two standard benchmarks: CIFAR-10 and ImageNet. In the case of CIFAR-10, we trained a small network of 3 convolutions followed by 2 fully connected layers. To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added. In the case of ImageNet, we used PyTorch’s pretrained VGG16 model for its convolution weights (which were frozen during training). We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization. We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImagnNet.
Figure 17: results for CIFAR-10 experiments. Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts; for instance, at epochs 3 and 4, which per the loss curves’ characteristics are where one would likely stop training. While the loss curves are similar in shape, we report that at epoch 4, for example, our training and validation losses are 25.6% and 5.1% lower than their standard network counterparts.
Figure 18: In our ImageNet test we similarly see modest gains of our (“experimental”) network over a Kaiming-initialized (“ordinary”) network, yet due to using a smaller learning rate here ($0.00001$, vs.0.0010.0010.0010.001 as used with CIFAR-10), the loss curves are smoother, so we see our networks consistently outperforming the standard network 25% into epoch 1. We trained both networks for 2 epochs; at epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).

### v6

@@ 4.2 Numerical Results
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization (See also Figure 5). Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks. Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
@@
Figure 4: Standard Kaiming initialization/gradient descent vs. pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
@@
Figure 5: Neuron outputs of a default (Kaiming-initialized) network (left) versus a pretrained variant of our network (right). The reason our network succeeds in using so many linear regions, when a random one fails, is that the number of regions is not a “local” property of parameter space. The first 3 layers of the default network are not adding any new linear regions, and infinitesimal adjustments to the slopes or biases of the lines shown would not change that. In our network, the pretraining guides it to a region of parameter space where $2^{d}$ activation regions are generated. Since the region count is not a local property, we still have a neighborhood of parameter space we can optimize in and maintain efficient neuron use.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$. Notice though that the individual neuron outputs are much messier than in our convex case (see Figure 5) This is because naive gradient descent has all the weights in between the two pretrained component networks to learn unassisted.
@@
The results presented so far are limited to one-dimensional convex functions, but our method can easily be extended. Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 8 in the Appendix for a schematic). For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs. Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
@@
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising. In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world regression datasets with up to eight input dimensions, as well as image classification problems.
@@
We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1. This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions. But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees. Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
@@
Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
@@ Appendix A Appendix
A.1 Initialization and Pretraining Algorithm
@@
     A-Gradient $\leftarrow\text{Derivative}(\text{Network},A)$Backpropagate through weight setting}
@@
     $A\leftarrow A-\epsilon\times\text{Gradient}$Update A, Not the network}
@@ A.3 Mathematical Results
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. This odd split around layer $i$ makes the proofs cleaner. Figure 9 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next. Valleys in one layer remain valleys in all future layers since 0 is a fixed point of each $T_{i}$. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition. On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected. Due to the reflection, the slopes of $W_{i}$ on each side of a peak or valley are proportional. Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$. Either way, it’s important to note that the slopes on each side of a bend scale identically during each composition.
@@ A.10 Real-World Data and Classification Problems
Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 14 and 15 already show impressive results using our current approach. In the case of the concrete problem (Figure 14), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced. In the case of the two-spiral classification task (Figure 15), our network is able to learn an accurate decision boundary (left subfigure), whereas a standard network constrained to be block diagonal fails to learn (right subfigure). When dense weight matrices are trained, our networks are almost able to tie with the standard initializations, although convergence can take longer. These results suggest that our method might be quite powerful when its parameter budget is spent on greater width and depth of convex function blocks (as in Figure 8) instead of filling in dense matrices.
@@
A.11 Larger Image Classification Examples
@@
We evaluated two standard benchmarks: CIFAR-10 and ImageNet. In the case of CIFAR-10, we trained a small network of three convolutions followed by two fully connected layers. To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added. In the case of ImageNet, we used PyTorch’s pretrained VGG16 model for its convolution weights (which were frozen during training). We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization. We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImageNet. Note that in both experiments, we switch parameterizations in our method halfway through training.
Figure 16: results for CIFAR-10 and ImageNet experiments. Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts. Interestingly, the results on ImageNet are much stronger (note the color labeling is the same for both images, our train and validation losses are the bottom two in ImageNet, whereas in Cifar-10 the two training losses are the lowest). This may be because the ImageNet experiments re-used the convolutional layers from VGG, rather than learning them on its own. It could also be in part due to larger layer sizes, since most activation functions are universal with infinite width, any detrimental effect of constraining them into convex shapes may be softened. In our ImageNet experiment we trained both networks for 2 epochs. Our method required a lower learning rate ($0.00001$) than used in previous experiments. At epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).